# Asymptotic stability of composite waves of two viscous shocks for relaxed compressible Navier-Stokes equations

### v2

@@ 4.3. Dissipative estimates
$\leq\frac{\mu}{4C_{1}}\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Phi\right)^{2}\mathrm{d}x\mathrm{d}t+C\tau^{2}\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}\mathrm{d}t+C\max\{\delta_{1},\delta_{2}\}.$
@@
$N_{1}^{k}$ $\leq\nu_{1}\tau\int_{\mathbb{R}}\left(\partial_{x}^{k}Q(t)\right)\mathrm{d}x+C(\nu_{1})\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi(t)\right)\mathrm{d}x+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t+C\tau\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t$
$+C\tau\int_{\mathbb{R}}\left(\partial_{x}^{k}Q_{0}\right)\mathrm{d}x+C\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi_{0}\right)\mathrm{d}x+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}\mathrm{d}t+C\max\{\delta_{1},\delta_{2}\}.$
@@
For $N_{3}^{k}$, using Young’s inequality, we have
$N_{3}^{k}\leq\frac{\mu}{8}\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k}Q\right)^{2}\mathrm{d}x\mathrm{d}t.$
@@
Proof of Proposition 3.1: Firstly, multiplying (4.16) and (4.17) by $1+2C_{2}$, $\frac{3}{2}C_{1}+2C_{1}C_{2}$, respectively, we have
$(\frac{1}{4}+C_{2})\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+\frac{1}{2}C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t\leq\nu(1+2C_{2})\|\Psi\|_{H^{1}}^{2}+\nu_{1}(\frac{3}{2}C_{1}+2C_{1}C_{2})\tau\|Q\|_{H^{1}}^{2}$ (4.18)
$\qquad+C(\nu_{1})\|\partial_{x}\Psi\|_{H^{1}}^{2}+C(\nu)\|\partial_{x}\Phi\|_{H^{1}}^{2}+C\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}|\dot{X}_{i}(t)|^{2}\mathrm{d}t+C_{3}\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t$
$\qquad+C\left(\|\Psi_{0}\|_{H^{2}}^{2}+\|\partial_{x}\Phi_{0}\|_{H^{1}}^{2}+\tau\|Q_{0}\|_{H^{1}}^{2}\right)+C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{0}^{t}(G^{s}(U)+G(U))\mathrm{d}t+C\delta_{0},$
where $C_{3}$ is a constant independent of $\tau$ and $T$. Then, by using the smallness of $\delta_{1},\delta_{2}$ and choosing $\nu,\nu_{1}$ to be sufficiently small in equation (4.18), we get
$(\frac{1}{4}+C_{2})\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+\frac{1}{2}C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t\leq\frac{1}{2}\|\Psi\|_{L^{2}}^{2}+\frac{1}{2}\tau\|Q\|_{L^{2}}^{2}$ (4.19)
$\qquad+C\left(\|(\partial_{x}\Psi,\partial_{x}\Phi,\sqrt{\tau}\partial_{x}Q)\|_{H^{1}}^{2}+\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t\right)+\frac{1}{16}\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}(t)|^{2}\mathrm{d}t$
$\qquad+C\left(\|\Psi_{0}\|_{H^{2}}^{2}+\|\partial_{x}\Phi_{0}\|_{H^{1}}^{2}+\tau\|Q_{0}\|_{H^{1}}^{2}\right)+\frac{1}{32}\int_{0}^{t}(G^{s}(U)+G(U))\mathrm{d}t+C\delta_{0}.$
Next, using Lemma 4.6 and choosing $\varepsilon_{1},\delta_{1},\delta_{2}$ sufficiently small, we get that
$(\frac{1}{8}+C_{2})\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+\frac{1}{4}C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t\leq\frac{1}{2}\|\Psi\|_{L^{2}}^{2}+\frac{1}{2}\tau\|Q\|_{L^{2}}^{2}+\frac{1}{2}\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}(t)|^{2}\mathrm{d}t$ (4.20)
$\qquad+\frac{1}{2}\int_{0}^{t}G(U)\mathrm{d}t+\frac{1}{2}\int_{0}^{t}G^{s}(U)\mathrm{d}t+C\left(\|\Psi_{0}\|_{H^{2}}^{2}+\|\partial_{x}\Phi_{0}\|_{H^{1}}^{2}+\tau\|Q_{0}\|_{H^{2}}^{2}\right)+C\delta_{0}.$
Then, combining the $L^{2}$ estimate (4.3) and (4.20), we get
$\|\Phi\|^{2}_{L^{2}}+\frac{1}{2}\|\Psi\|^{2}_{L^{2}}+\frac{1}{2}\tau\|Q\|^{2}_{L^{2}}+\frac{1}{2}\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}\mathrm{d}t+\frac{1}{2}\int_{0}^{t}\left(G^{s}(U)+G(U)\right)\mathrm{d}t$ (4.21)
$\qquad\qquad+\frac{1}{8}\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+\frac{1}{4}C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t$
$\leq C\left(\|\Psi_{0}\|_{H^{2}}^{2}+\|\Phi_{0}\|_{H^{2}}^{2}+\tau\|Q_{0}\|_{H^{2}}^{2}\right)+C\delta_{0}.$
Thus, combing the high-energy estimates (4.13) and choosing $\varepsilon_{1},\delta_{1},\delta_{2}$ sufficiently small, we derive that
$\|\Phi\|^{2}_{H^{2}}+\|\Psi\|^{2}_{H^{2}}+\tau\|Q\|^{2}_{H^{2}}+\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}\mathrm{d}t+\int_{0}^{t}\left(G^{s}(U)+\|\left(\Phi_{x},\Psi_{x}\right)\|^{2}_{H^{1}}+\|Q\|^{2}_{H^{2}}\right)\mathrm{d}t$ (4.22)
$\leq C\left(\|\Phi_{0}(\cdot)\|^{2}_{H^{2}}+\|\Psi_{0}(\cdot)\|^{2}_{H^{2}}+\tau\|Q_{0}(\cdot)\|^{2}_{H^{2}}\right)+C\delta_{0}.$
This finish the proof of Proposition 3.1.
@@
$(\Phi^{\tau},\Psi^{\tau})\rightharpoonup(\Phi^{0},\Psi^{0})\qquad weakly-*\quad in\quad L^{\infty}((0,\infty);H^{2}),$ (5.1)

### v3

@@ 4.3. Dissipative estimates
$\tau\int_{\mathbb{R}}\partial_{x}^{k}Q(t)\partial_{x}^{k+1}\Psi(t)\mathrm{d}x\leq\nu_{1}\tau\int_{\mathbb{R}}(\partial_{x}^{k}Q(t))^{2}\mathrm{d}x+C(\nu_{1})\tau\int_{\mathbb{R}}(\partial_{x}^{k+1}\Psi(t))^{2}\mathrm{d}x,$
$\tau\int_{\mathbb{R}}\int_{\mathbb{R}}\partial_{x}^{k}Q_{0}\partial_{x}^{k+1}\Psi_{0}\mathrm{d}x\leq C\tau\int_{\mathbb{R}}(\partial_{x}^{k}Q_{0})^{2}\mathrm{d}x+C\tau\int_{\mathbb{R}}(\partial_{x}^{k+1}\Psi_{0})^{2}\mathrm{d}x,$
@@
$\leq\mu\kappa\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Phi\right)^{2}\mathrm{d}x\mathrm{d}t+C\tau^{2}\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}\mathrm{d}t+C\max\{\delta_{1},\delta_{2}\}.$
@@
$N_{1}^{k}$ $\leq\nu_{1}\tau\int_{\mathbb{R}}\left(\partial_{x}^{k}Q(t)\right)^{2}\mathrm{d}x+C(\nu_{1})\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi(t)\right)^{2}\mathrm{d}x+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t+C\tau\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t$
$+C\tau\int_{\mathbb{R}}\left(\partial_{x}^{k}Q_{0}\right)^{2}\mathrm{d}x+C\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi_{0}\right)^{2}\mathrm{d}x+C\delta_{0}\int_{0}^{t}G^{s}\mathrm{d}t+\mu\kappa\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Phi\right)^{2}\mathrm{d}x\mathrm{d}t+C\max\{\delta_{1},\delta_{2}\}.$
@@
For $N_{3}^{k}$, using Young’s inequality and noting that $|v_{m}-v_{-}|\leq C\delta_{1},|v_{m}-v_{+}|\leq C\delta_{2}$, we have
$N_{3}^{k}\leq\frac{\mu}{2}\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t+\frac{v_{m}+C(\delta_{0}+\varepsilon_{1})}{2}\int_{0}^{t}\int_{\mathbb{R}}\frac{v}{\mu}\left(\partial_{x}^{k}Q\right)^{2}\mathrm{d}x\mathrm{d}t.$
@@
Proof of Proposition 3.1: For arbitrary constants $C_{2},C_{3},C_{4}>0$, we multiply equations (4.3), (4.18) and (4.19) by $C_{2}$, $\frac{3\mu}{4v_{m}}C_{3}$ and $\frac{2}{v_{m}}C_{4}$, respectively. Combining the resulting equations with Lemma 4.7 yields
$\frac{1}{2}\|\Psi\|^{2}_{L^{2}}+\int_{\mathbb{R}}H(v|\widetilde{v})\mathrm{d}x+\frac{\tau}{2\mu}\|Q\|^{2}_{L^{2}}+\|\partial_{x}\Phi\|^{2}_{H^{1}}+\|\partial_{x}\Psi\|^{2}_{H^{1}}+\tau\|\partial_{x}Q\|^{2}_{H^{1}}$ (4.20)
$+\sum\limits_{i=1}\limits^{2}(\frac{\delta_{i}}{4M}-C\delta_{i}^{2})\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t+C_{1}(1-C(\delta_{0}+\varepsilon_{1}))\int_{0}^{t}G^{s}(U)\mathrm{d}t+C_{5}\int_{0}^{t}G(U)\mathrm{d}t$
$+C_{6}\int_{0}^{t}\|\sqrt{|p^{\prime}(v)|}\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+C_{7}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t+\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t$
$\leq C\left(\|\Phi_{0}(\cdot)\|^{2}_{H^{2}}+\|\Psi_{0}(\cdot)\|^{2}_{H^{2}}+\tau\|Q_{0}(\cdot)\|^{2}_{H^{2}}\right)+\nu\frac{3\mu}{4v_{m}}C_{3}\|\Psi\|_{L^{2}}^{2}+\nu_{1}\frac{2C_{4}}{v_{m}}\tau\|Q\|_{L^{2}}^{2},$
where
$C_{5}=C_{2}(1-\kappa-C(\delta_{1}^{\frac{1}{2}}+\delta_{2}^{\frac{1}{2}}))-C_{4}(1+C(\delta_{0}+\varepsilon_{1})),$
$C_{6}=\frac{3\mu}{4v_{m}}\left(C_{3}(1-\kappa-C(\delta_{0}+\varepsilon_{1}))-C_{2}\frac{1+\kappa+C\varepsilon_{1}}{1-C\varepsilon_{1}}\right),$
$C_{7}=\frac{\mu}{v_{m}}\left(C_{4}(1-C\delta_{0})-\frac{3}{4}C_{3}(1+C\delta_{0})\right).$
We now choosing $\delta_{1},\delta_{2},\delta_{0},\varepsilon_{1},\kappa,\nu_{1},\nu_{2}$ sufficiently small such that
$\frac{1}{4M}>C\delta_{i},\quad 1>C(\varepsilon_{1}+\delta_{0}),\quad\nu_{1}\frac{3\mu C_{3}}{4v_{m}}<\frac{1}{2},\quad\nu_{2}\frac{2C_{4}}{v_{m}}<\frac{1}{2\mu}\quad\text{where}\quad i=1,2.$
Furthermore, to assure $C_{5},C_{6},C_{7}>0$, it is sufficient to require $C_{3}>C_{2}>C_{4}>\frac{3}{4}C_{3}$. Therefore, combining the above results and using Lemma 4.1, the proof of the Proposition 3.1 is completed.
@@
$(\Phi^{\tau},\Psi^{\tau})\rightharpoonup(\Phi^{0},\Psi^{0})\qquad weak-*\quad in\quad L^{\infty}((0,\infty);H^{2}),$ (5.1)